OpenStudy (anonymous):

This is a Quadratic Equation, solve it by factoring please.Also please explain. 3x^2=-4x-1

OpenStudy (anonymous):

3x(2)=-4x-1 Multiply 3x by 2 in the numerator. Multiply 3x by 2to get 6x. 6x=-4x-1 Since -4x contains the variable to solve for, move it to the left-hand side of the equation by adding 4x to both sides. The -4x will cancel out on the right-hand side and look like this. 6x+4x=-4x+4x-1 6x+4x=-1 Since 6x and 4x are like terms, add 4x to 6x to get 10x. 10x=-1 Divide each term in the equation by 10. 10x/10=-1/10 Simplify the left-hand side of the equation by canceling the common factors. x= -1/10

OpenStudy (anonymous):

Thank you tons, and great explanation